If x= 1 ____ , find the value x 3 - 2x²-7x+5 2- √3


If x=    1


         ____ , find the value  x3 - 2x²-7x+5


         2-√3


 


Grade:9

2 Answers

Godfrey Classic Prince
12 years ago

Dear ARTI SHARMA,

I cannot CLEARLY UNDERSTAND IT..

Please upload the problem and I will help you for sure..

 

All the Very Best & Good Luck to you ..

 

Best Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-M

 

pankaj kumar raman
12 years ago

x=1/2-√3

=> 2x-√3x=1

=> 2x-1=√3x

squaring on both the sides

=> 4x² + 1-4x = 3x²

=> x²-4x+1=0=q(x)...[say]....................(1)

 

divide this equation by x³ - 2x²-7x+5=p(x)....[say]

=>p(x)=q(x).(x+2)+3..                                                  [please solve it]

          from(1)

            =0.(x+2)+3

            =3 ans.